# ResGS: Residual Densification of 3D Gaussian for Efficient Detail Recovery

### v1

@@ 4.2 Comparisons
To quantitatively assess the visual quality of our method, we compared with several 3D-GS variants: 3D-GS [26], Scaffold-GS [31], Octree-GS [39], FreGS [54], AbsGS [49], Pixel-GS [56] and Mini-Splatting-D [13]. Note that Scaffold-GS and Octree-GS use a latent representation rather than the original 3D-GS representation. Some radiance-field-based methods are also selected for comparison, including Plenoxels [51], Instant-NGP [34] and Mip-NeRF360 [1]. Note that the resolution of Mip-NeRF360 dataset images varies across different works. Following the resolution setup in the 3D-GS work, we retrain some models on the Mip-NeRF360 dataset. Additionally, for works that did not report their storage, we obtained and reported it based on their released code. As shown in Table 1, our method achieves SOTA performance across all metrics on the Mip-NeRF360 dataset. On the Tanks&Temples dataset, we also achieve the best SSIM and LPIPS metrics, which closely reflect human perception. In the Deep Blending dataset, our method performs comparably to the best results of existing approaches. Additionally, our model variant without AbsGS [49] maintains similar performance to the full model, with only a slight reduction in LPIPS. Notably, our smaller model version achieves strong performance with reduced memory usage, showcasing the efficiency and scalability of our approach.
@@
Table 2: Analyzing residual split upon multiple 3D-GS variants. The performance consistently increases on these pipelines, indicating that our Residual split outperforms the split and clone strategy using a fixed threshold.
@@
In Fig. 5, we present qualitative comparisons between our method and several 3D-GS variants [26, 56, 49] across various scenes, including both indoor and outdoor settings. It is evident that our model captures finer details, as shown in the second, third, and fourth rows, constructs missing geometry more effectively (last row), and achieves more accurate geometry reconstruction (first row).
@@ 4.3.1 Residual Split
Analyzing $\beta$.
To analyze the impact of different opacity reduction weights $\beta$ on rendering quality, we tested our model upon the Tanks&Temples dataset while applying different $\beta$. All the models are trained with similar storage. The results are shown in Fig. 4. While $\beta$ is sampled in the range of $(0.05,1.0)$, the LPIPS scores do not differ much when $\beta<0.4$. A notable performance drop (increasing LPIPS) is observed.
@@
We analyze the compatibility of our novel densification method by replacing split and clone in multiple 3D-GS variants [26, 49, 56, 13] with residual split. The hyper-parameters were kept the same as in the original pipelines. The results are shown in Table 2. The table shows that our method leads to a notable increase in PSNR metrics across all pipelines, while for other metrics, some exhibit improvements, and others remain approximately the same. Additionally, memory consumption is reduced after incorporating residual split compared to the original pipeline. This underscores our method’s ability to improve rendering quality and reduce redundancy, while also demonstrating its compatibility and potential for broader application across 3D-GS-oriented works as an effective alternative to split and clone.
4.3.2 Split-Stage Schedule
Figure 6: Impact of the split-stage schedule. We show the Gaussian ellipsoids and the rendered images at the end of Stage 1, Stage 2, and the final stage. Our model constructs a coarse structure in the early stages and refines details in the later stages.
To evaluate the functionality of our split-stage training schedule, visualizations of intermediate results are presented in Fig. 6. At the early stages, due to the absence of high-frequency information, our model captures the overall structure of the scene, using a few large, coarse Gaussians to represent it. As training progresses to later stages, more intricate information is revealed, allowing finer details to be carved out.
@@ 4.3.3 Ablation Studies
To assess the effectiveness of individual modules in our proposed pipeline—residual split, image pyramid as supervision, and varying gradient threshold—ablation studies on the Mip-NeRF360 [1] dataset were conducted. For the varying gradient threshold, we only conducted studies on its removal, as it is highly interdependent with the other two modules. The results are shown in Table 3. It is evident that both residual split and the image pyramid supervision individually enhance the quality of our baseline, and combining them results in even better performance. Additionally, applying the varying gradient threshold further boosts performance. In conclusion, each module contributes to improved results.
@@ 4.4 Discussions and Limitations
Our results show that our method significantly improves rendering quality on the Mip-NeRF360 [1] and Tanks&Temples [27] datasets. However, for Deep Blending [20], our method did not yield a significant boost in performance. This variation is primarily due to the challenges specific to each dataset: Mip-NeRF360 and Tanks&Temples require capturing fine textures and recovering missing geometry, whereas Deep Blending on managing occlusions and preventing overfitting to training views. Our method mainly contributes to retrieving fine details and geometry and does not address the issue of overfitting and occlusion, suggesting it to be a topic for future optimization.
@@ 5 Conclusion
In this paper, we analyzed the limitations of the current 3D-GS densification method. To address these issues, we propose residual split, a residual empowered densification method that adaptively handles various cases and fits the scene progressively. Additionally, we introduce ResGS, a coarse-to-fine 3D-GS pipeline to supplement residual split. Specifically, we use an image pyramid for supervision to gradually recover the overall structure to fine details during Gaussian optimization. We also design a Gaussian selection scheme to further densify coarse Gaussians progressively. Experiments show our pipeline achieves SOTA rendering quality while residual split is highly compatible with other 3D-GS variants.
@@
6 More Experiment Results
In this section, we give more experimental results of our work.
Variance of PSNR on Deep Blending Dataset.
@@ 6 More Experiment Results
Further Analyzing Residual Split on 3D-GS.
Periodic Opacity Reduction.
Per Scene Results.
We present the per-scene results of the used metrics in Table 9-14. For works that did not report per-scene results, we obtained and reported them using their released code. The works selected for comparison are mainly the same as in the paper: 3D-GS [26], Scaffold-GS [31], Octree-GS [39], AbsGS [49], Pixel-GS [56], Mini-Splatting-D [13], Plenoxels [51], Instant-NGP [34] and Mip-NeRF360 [1]. Note that we did not show the results of FreGS [54] since they did not report their per-scene results or release their code. Works retrained on the Mip-NeRF360 [1] are marked with a .
Image Resolution of Mip-NeRF360 Dataset.
In our paper, we mentioned that the image resolution on the Mip-NeRF360 [1] dataset varies across different works. 3D-GS [26] uses the provided “images_4”, the official downscaled 4 times images from the dataset, for outdoor scenes and “images_2” for indoor scenes. Most works [49, 13, 54] follow the 3D-GS setting, but Scaffold-GS [31], Octree-GS [39], Pixel-GS [56] uses different settings. Specifically, Scaffold-GS and Octree-GS downscale original images to 1.6K resolution, while Pixel-GS downscales original images by a factor of 4 for outdoor scenes and 2 for indoor scenes during the training process, instead of using the provided downscaled images. This discrepancy results in differences in experimental settings. To further demonstrate the performance of our method, we provide its results under the two different settings, as shown in Table 15 and 16. The tables show that our method prevails even under these alternate settings.
@@
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Scaffold-GS 27.71 0.813 0.221
Octree-GS 27.73 0.815 0.217
Ours-Small (AbsGS) 28.04 0.831 0.191
Ours (3D-GS) 28.10 0.832 0.187
Ours (AbsGS) 28.14 0.834 0.174
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Pixel-GS 27.88 0.834 0.176
Ours-Small (AbsGS) 28.21 0.841 0.176
Ours (3D-GS) 28.28 0.843 0.172
Ours (AbsGS) 28.30 0.845 0.160

### v2

@@ 4.2 Comparisons
To assess the effectiveness of our method, we compared it with several 3D-GS variants: 3D-GS [24], Scaffold-GS [29], Octree-GS [35], FreGS [48], AbsGS [43], Pixel-GS [50] and Mini-Splatting-D [13]. Note that Scaffold-GS and Octree-GS use a latent representation rather than the original 3D-GS representation. Some radiance-field-based methods are also selected for comparison, including Plenoxels [45], Instant-NGP [32] and Mip-NeRF360 [1]. We found that the resolution of Mip-NeRF360 dataset images varies across different works. Following the resolution setup in the 3D-GS work, we retrain some models on the Mip-NeRF360 dataset. Additionally, for works that did not report their storage, we obtained and reported it based on their released code.
Quality Comparisons.
In Table 1, we present quantitative results comparing the rendering quality of our method with other approaches. From the table, we can see that our method achieves SOTA performance across all metrics on the Mip-NeRF360 [1] dataset. On the Tanks&Temples [25] dataset, we also achieve the best SSIM and LPIPS metrics, which closely reflect human perception. In the Deep Blending [18] dataset, while our method does not achieve the highest score, it shows a notable improvement over the baseline. Additionally, our model variant without AbsGS [43] maintains similar performance to the full model, with only a slight reduction in LPIPS. Notably, our smaller model version achieves strong performance with reduced memory usage, showcasing the efficiency and scalability of our approach. In Fig. 5, we present qualitative comparisons between our method and several 3D-GS variants [24, 50, 43] across various scenes, including both indoor and outdoor settings. It is evident that our model captures finer details, as shown in the second, third, and fourth rows, constructs missing geometry more effectively (last row), and achieves more accurate geometry reconstruction (first row).
@@
Table 2: Efficiency comparison with former methods.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
FPS Time FPS Time FPS Time
Octree-GS 119 29m55s 126 23m8s 151 34m38s
Scaffold-GS 103 23m53s 137 15m37s 141 19m40s
3D-GS 103 27m16s 127 15m34s 101 24m48s
AbsGS 132 28m48s 193 13m35s 185 22m3s
Mini-Splatting-D 74 34m42s 75 27m39s 100 30m37s
Pixel-GS 65 40m6s 70 29m29s 70 34m1s
Ours-Small (AbsGS) 141 20m56s 206 11m35s 202 19m26s
Ours (AbsGS) 100 29m24s 150 16m29s 153 23m33s
Table 3: Analyzing residual split upon multiple 3D-GS variants. The performance consistently increases on these pipelines, indicating that our Residual split is a better densification operation. 3D-GS indicates a retrained version of 3D-GS [24] for better performance.
@@
Efficiency Comparisons.
Here, we provide the computational costs of our method in comparison with other approaches. The results are shown in Table. 2. From the table, we can see that our method involves a slight increase in training time while delivering much higher rendering quality than methods with lower rendering quality like 3D-GS [24] and AbsGS [43]. Our method is also significantly more efficient than other methods like Mini-Splatting-D [13], with comparable rendering quality. Notably, our small model provides the fastest FPS and shortest training time in all compared methods while maintaining high fidelity.
@@ 4.3.1 Residual Split
We analyze the compatibility of our novel densification operation by replacing split and clone in multiple 3D-GS variants [24, 43, 50, 13] with residual split. The hyperparameters were kept the same as in the original pipelines. The results are shown in Table 3. The table shows that our method leads to a notable increase in PSNR metrics across all pipelines, while for other metrics, some exhibit improvements, and others remain approximately the same. Additionally, memory consumption is reduced after incorporating residual split compared to the original pipeline. This underscores our method’s ability to improve rendering quality and reduce redundancy, while also demonstrating its compatibility and potential for broader application across 3D-GS-oriented works as an effective alternative to split and clone.
Figure 7: Impact of the scale reduction factor $\lambda_{s}$ for residual split.
Hyperparameter Sensitivities.
Since residual split involves some hyperparameters: opacity reduction weight $\beta$ and scale reduction factor $\lambda_{s}$, we conducted a corresponding sensitivity analysis. We tested our method upon the Tanks&Temples dataset while applying different $\beta\in[0.05,1.0)$ and $\lambda_{s}\in[1.05,2.0]$. All the models are trained with similar storage. The LPIPS scores are shown in Fig. 6 and 7. As shown in the figures, LPIPS scores remain relatively stable when $\beta<0.4$ and $\lambda_{s}>1.55$. These results indicate that our method is robust to hyperparameter variations, with only a 0.006 difference in LPIPS between the best and worst cases, and a relatively broad range where rendering quality remains optimal.
@@ 4.3.2 Ablation Studies
To assess the effectiveness of individual modules in our proposed pipeline—residual split, image pyramid as supervision, and varying gradient threshold—ablation studies on the Mip-NeRF360 [1] dataset were conducted. For the varying gradient threshold, we only conducted studies on its removal, as it is highly interdependent with the other two modules. The results are shown in Table 4. It is evident that both residual split and the image pyramid supervision individually enhance the quality of our baseline. However, we observe that adding image pyramid supervision alone does not lead to significant improvements. This is because the split-and-clone operations do not perform well within a coarse-to-fine schedule. In the early stages, when the overall structure is still being constructed, split is primarily chosen, which exacerbates the issue rather than improving it. On the other hand, in later iterations, when fine details need to be recovered, clone is selected and does not address the issue properly. Conversely, residual split works perfectly due to its adaptiveness. This is evident from the notable improvement of Base + RS + IP over both Base + IP and Base + RS, which also shows that a coarse-to-fine schedule supplements residual split as expected. Additionally, applying the varying gradient threshold further boosts performance. In conclusion, each module contributes to improved results.
@@ 4.4 Discussions and Limitations
Our results show that our method significantly improves rendering quality on the Mip-NeRF360 [1] and Tanks&Temples [25] datasets. However, for Deep Blending [18], our method did not yield a significant boost in performance. This variation is primarily due to the challenges specific to each dataset: Mip-NeRF360 and Tanks&Temples require capturing fine textures and recovering missing geometry, whereas Deep Blending focuses on handling occlusions and reconstructing accurate geometry in weakly textured areas. Our method mainly contributes to retrieving fine details and sufficient geometry and does not address the issue of occlusion and weak-textured areas, suggesting it to be a topic for future optimization.
@@ 5 Conclusion
In this paper, we analyzed the limitations of the current 3D-GS densification operation. To address these issues, we propose residual split, a residual empowered densification operation that adaptively handles both over-reconstruction and under-reconstruction. Additionally, we introduce ResGS, a coarse-to-fine 3D-GS pipeline to supplement residual split. Specifically, we use an image pyramid for supervision to gradually recover the overall structure to fine details during Gaussian optimization. We also design a Gaussian selection scheme to further densify coarse Gaussians progressively. Experiments show our pipeline achieves SOTA rendering quality while residual split is highly compatible with other 3D-GS variants.
@@
6 Additional Experiment Results
In this section, we provide additional experimental results of our work.
6.1 Variance of PSNR on Deep Blending Dataset
@@ 6.2 Further Analysis of Residual Split on 3D-GS
6.3 Periodic Opacity Reduction
6.4 Image Resolution of Mip-NeRF360 Dataset
In our paper, we mentioned that the image resolution on the Mip-NeRF360 [1] dataset varies across different works. 3D-GS [24] uses the provided “images_4”, the official downscaled 4 times images from the dataset, for outdoor scenes and “images_2” for indoor scenes. Most works [43, 13, 48] follow the 3D-GS setting, but Scaffold-GS [29], Octree-GS [35], Pixel-GS [50] uses different settings. Specifically, Scaffold-GS and Octree-GS downscale original images to 1.6K resolution, while Pixel-GS downscales original images by a factor of 4 for outdoor scenes and 2 for indoor scenes during the training process, instead of using the provided downscaled images. This discrepancy results in differences in experimental settings. To further demonstrate the performance of our method, we provide its results under the two different settings, as shown in Table 10 and 11. The tables show that our method prevails even under these alternate settings.
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Scaffold-GS 27.71 0.813 0.221
Octree-GS 27.73 0.815 0.217
Ours-Small (AbsGS) 28.04 0.831 0.191
Ours (3D-GS) 28.10 0.832 0.187
Ours (AbsGS) 28.14 0.834 0.174
Dataset Mip-NeRF360
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Pixel-GS 27.88 0.834 0.176
Ours-Small (AbsGS) 28.21 0.841 0.176
Ours (3D-GS) 28.28 0.843 0.172
Ours (AbsGS) 28.30 0.845 0.160
6.5 Per Scene Results
We present the per-scene results of the used metrics in Table 12-17. For works that did not report per-scene results, we obtained and reported them using their released code. The works selected for comparison are mainly the same as in the paper: 3D-GS [24], Scaffold-GS [29], Octree-GS [35], AbsGS [43], Pixel-GS [50], Mini-Splatting-D [13], Plenoxels [45], Instant-NGP [32] and Mip-NeRF360 [1]. Note that we did not show the results of FreGS [48] since they did not report their per-scene results or release their code. Works retrained on the Mip-NeRF360 [1] are marked with a .